(D4 - 64) y= x cos x


1
Expert's answer
2021-06-02T14:40:27-0400

"\left( {{D^4} - 64} \right)y = x\cos x \Rightarrow {y^{(4)}} - 64y = x\cos x"

The characteristic equation has the form

"{k^4} - 64 = 0"

"({k^2} - 8)({k^2} + 8) = 0"

"(k - \sqrt 8 )(k + \sqrt 8 )(k - \sqrt 8 i)(k + \sqrt 8 i) = 0"

"{k_{1,2}} = \pm \sqrt 8 ,\,\,{k_{3,4}} = \pm \sqrt 8 i"

Then the general solution of the homogeneous equation is

"{y_h} = {C_1}{e^{\sqrt 8 x}} + {C_2}{e^{ - \sqrt 8 x}} + {C_3}\cos \sqrt 8 x + {C_4}\sin \sqrt 8 x"

We will seek a particular solution in the form

"Y = (Ax + B)\cos x + (Cx + D)\sin x \Rightarrow Y' = A\cos x - (Ax + B)\sin x + C\sin x + (Cx + D)\cos x = (Cx + D + A)\cos x + ( - Ax - B + C)\sin x \Rightarrow"

"\Rightarrow Y'' = C\cos x - (Cx + D + A)\sin x - A\sin x + ( - Ax - B + C)\cos x = ( - Ax - B + 2C)\cos x + ( - Cx - D - 2A)\sin x \Rightarrow"

"\Rightarrow Y''' = - A\cos x - ( - Ax - B + 2C)\sin x - C\sin x + ( - Cx - D - 2A)\cos x = ( - Cx - D - 3A)\cos x + (Ax + B - 3C)\sin x \Rightarrow"

"\Rightarrow {Y^{(4)}} = - C\cos x - ( - Cx - D - 3A)\sin x + A\sin x + (Ax + B - 3C)\cos x = (Ax + B - 4C)\cos x + (Cx + D + 4A)\sin x"

Substitute into the equation:

"(Ax + B - 4C)\cos x + (Cx + D + 4A)\sin x - 64\left( {(Ax + B)\cos x + (Cx + D)\sin x} \right) = x\cos x"

"x\cos x(A - 64A) + \cos x(B - 4C - 64B) + x\sin x(C - 64C) + \sin x(D + 4A - 64D) = x\cos x"

"\left\{ \begin{array}{l}
 - 63A = 1\\
 - 63B - 4C = 0\\
 - 63C = 0\\
4A - 63D = 0
\end{array} \right."

"A = - \frac{1}{{63}},\,\,B = 0,\,C = 0,\,D = - \frac{4}{{3969}}"

Then

"Y = - \frac{1}{{63}}x\cos x - \frac{4}{{3969}}\sin x"

"y = {y_h} + Y = {C_1}{e^{\sqrt 8 x}} + {C_2}{e^{ - \sqrt 8 x}} + {C_3}\cos \sqrt 8 x + {C_4}\sin \sqrt 8 x - \frac{1}{{63}}x\cos x - \frac{4}{{3969}}\sin x"

Answer: "y = {C_1}{e^{\sqrt 8 x}} + {C_2}{e^{ - \sqrt 8 x}} + {C_3}\cos \sqrt 8 x + {C_4}\sin \sqrt 8 x - \frac{1}{{63}}x\cos x - \frac{4}{{3969}}\sin x"